OpenStudy (anonymous):

Peter needs 12 quarts of paint to paint a room of his house. If 1 gallon is equal to 4 quarts, how many gallons of paint should Peter buy?

OpenStudy (anonymous):

will fan and medal

OpenStudy (anonymous):

3

OpenStudy (anonymous):

no guessing

OpenStudy (anonymous):

x/1=12/4 therefore he needs 3 gallons

OpenStudy (tkhunny):

Convert the units: \(12\;quarts\cdot\dfrac{1\;gallon}{4\;quarts} = \)?? What do you get?